Five students are equally qualified to represent their schoolat an awards ceremony. Two students will be chosen randomly to go to the ceremony.
A.) If Tony is one fo the five students,what is his probability of being chosen?
B.)Tony's friend, Allecia who was also eligible to go has been chosen. What is Tony's probabilityof being chosen when he learns Allecia is going?

Answer by stanbon(57290)

Five students are equally qualified to represent their schoolat an awards ceremony. Two students will be chosen randomly to go to the ceremony.
A.) If Tony is one fo the five students,what is his probability of being chosen?
# of pairs with tony in the pair: 1*4=4
---
3 of pairs without restriction: 5C2 = 10
---
P(Tony is in the pair chosen) = 4/10 = 2/5
======================================================
B.)Tony's friend, Allecia who was also eligible to go has been chosen. What is Tony's probability of being chosen when he learns Allecia is going?
P(tony chose | alecia chosen) = 1/4
=================================================
Cheers,
Stan H.